Exactly 1.5g of soil is saturated with 100 mL of 1 M KCl. At the end of the saturation, the solution is analyzed, and the following results obtained: Ca = 0.50 mg/L; Mg = 1 mg/L; Na = 1.5mg/L; Al = 2 mg/L. If this represent all of the exchangeable ions that were initially on the soil, what is the CEC and Base Saturation in men/100g?
Answer #1

1.5 g of soil is saturated with 100 ml of 1 M KCl. The analyzed concentrations are:

[Ca2+] = 0.50 mg/L, [Mg2+] = 1 mg/L, [Na+]=1.5 mg/L, [Al3+] = 2 mg/L.

Cation exchange capacity (CEC) is a measure of the total negative charges within the soil that adsorb plant nutrients cations such as Ca, Mg, Na ions, etc.

The milliequivalents can be calculated as follows:

[Ca2+]meq = 0.0005/20 g/L = 0.000025 g/L.

[Mg2+]meq = 0.001/12 g/L = 0.000083 g/L.

[Na+]meq = 0.0015/23 g/L = 0.000065 g/L.

[Al3+]meq = 0.002/9 g/L = 0.000222 g/L.

Total milliequivalents = [Ca2+]meq + [Mg2+]meq + [Na+]meq + [Al3+]meq

= 0.000025 + 0.000083 + 0.000065 + 0.000222 g/L = 0.000395 g/L.

So, 1.5 g of soil contains 0.395 ppm of cations.

100 g of soil will contain

So, the total CEC is 0.026 meq/100g of soil.

Base saturation is the percentage of the CEC occupied by the basic cations Ca2+, Mg2+, and Na+. Al3+ is an acid cation, so it is not considered in base saturation.

Base saturation

So, the base saturation is 43.94%.